OpenStudy (anonymous):

Solve for x: 3 over x minus 2 over 3 equals 5 over 6

OpenStudy (campbell_st):

is the equation \[\frac{3}{\frac{x -2}{3}} = \frac{5}{6}\]

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok... so looking at the left hand fraction, use the rule for dividing by fractions, find the reciprocal and multiply so the question becomes \[\frac{3}{1} \times \frac{3}{(x -2)} = \frac{5}{6}\] simplify this step then cross multiply to remove the denominators.. this would then allow you to then solve for x.

OpenStudy (anonymous):

Okay

OpenStudy (campbell_st):

The answer will be a mixed number or decimal... just depends on how you write it

OpenStudy (anonymous):

Okay Thanks